An electron and a proton are each placed at rest in a uniform electric field of magnitude 494 N/C. Calculate the speed of each particle 44.4 ns after being released. electron proton m/s m/s
Answer #1

here,

for electron

charge , q = 1.6 * 10^-19 C

mass , m = 9.1 * 10^-31 kg

the magnitude of electric feild , E = 494 N/C

accelration , a = q * E /m

a = 1.6 * 10^-19 * 494 /(9.1 * 10^-31)

a = 8.69 * 10^13 m/s^2

time interval , t = 44.4 ns = 4.44 * 10^-8 s

the speed of particle , v = 0 + a * t

v = 0 + 8.69 * 10^13 * 4.44 * 10^-8 m/s

v = 3.86 * 10^6 m/s

for proton

charge , q = 1.6 * 10^-19 C

mass , m = 1.67 * 10^-27 kg

the magnitude of electric feild , E = 494 N/C

accelration , a = q * E /m

a = 1.6 * 10^-19 * 494 /(1.67 * 10^-27)

a = 4.73 * 10^10 m/s^2

time interval , t = 44.4 ns = 4.44 * 10^-8 s

the speed of particle , v = 0 + a * t

v = 0 + 4.73 * 10^10 * 4.44 * 10^-8 m/s

v = 2.1 * 10^3 m/s